# Scaling Limit of a Stochastic Clustering Model on $\mathds{R}$

### v1

@@ Abstract
We consider an infinite-dimensional stochastic clustering model on $\mathds{R}$. In discrete time, each point of a unit-intensity simple point process moves halfway toward either of its left or right neighbors, chosen uniformly at random. Co-located points are merged into a single point, and the resulting simple point process is rescaled to unit intensity. We show that, when the point processes are shifted so that there is a point at the origin, there is a unique weak limit of these dynamics when the initial point process is renewal. We also conjecture that for the time-reversed process and with an appropriate scaling in space, there is a limiting (random) distribution function on $\mathds{R}$, whose corresponding measure assigns to $\mathds{R}$ a measure corresponding to the gap between consecutive points. Finally, we discuss some relevant research directions.
@@ 1 Introduction
Dynamic clustering is often done via numerical optimization [li2021survey]. For large, finite datasets, there are two challenges with this approach. First, large numerical optimization problems are computationally expensive. Second, and perhaps more important, is the question of determining stopping criteria for the clustering algorithm. For finite datasets, continued use of a clustering algorithm often results in all data points being assigned to a single cluster, which is undesirable.
In this work, we address this second question. Specifically, we study stochastic dynamic clustering algorithms on infinite datasets, aiming to determine whether these dynamics possess a stationary measure. If a stationary measure exists, a natural stopping criterion for the clustering dynamics is when one is close to that stationary measure. For large finite datasets, one can also accept the current clustering when the cluster gap distribution is close to that of the stationary measure.
To our knowledge, this is the first work to directly analyze this type of clustering dynamics on an infinite dataset, with the intent of informing behavior on large finite datasets. Indeed, we demonstrate this behavior via simulation for two simple algorithms below. These algorithms are as follows.
@@ 1.1 Related Work
Multiagent averaging dynamics have been studied extensively from a control perspective through various models with applications in distributed coordination and formation [hendrickx2012convergence, jadbabaie2003coordination], animal flocking [chazelle2014convergence], and opinion dynamics [degroot1974reaching, friedkin1999social, hegselmann2002opinion]. While averaging dynamics were initially used to model consensus in multiagent systems (e.g., robotic rendezvous [bullo2009distributed] or consensus in social networks [degroot1974reaching]), these models were later extended to study disagreement (clustering) in addition to consensus. For instance, in the context of opinion dynamics, Friedkin and Johnsen were the first to take this approach by proposing a model that captures opinion polarization into multiple clusters [friedkin1999social]. This model was later extended by Hegselmann and Krause [hegselmann2002opinion], and is known to belong to a broader class of bounded-confidence dynamics [bernardo2024bounded]. Unfortunately, analyzing the convergence and limit points of such dynamics poses major challenges, in contrast to standard averaging models, whose convergence properties can often be studied using classical tools from Markov chains or graph theory. These difficulties become even more pronounced when additional stochasticity is present in the evolution of the dynamics.
@@
It turns out that averaging dynamics on graphs that depend on agents’ states (e.g., agents’ positions) are closely related to dynamic clustering, especially when the data points are spatially located and evolve over time, and the goal is to cluster the points based on their similarities [aggarwal2003framework, Bhatti2019]. Similarly, in the context of dynamic community detection, one aims to explore how diffusion-based averaging dynamics on time-varying graphs can reveal communities (clusters) that persist over time [delvenne2010stability]. In both such applications, the co-evolution of agents’ states and the network structure introduces a dynamic coupling among the agents’ evolutions, which requires new analytical approaches to characterize the limit points and long-term behavior of these dynamics.
@@
Motivated by these applications, we propose and study a new class of averaging dynamics for spatial data clustering, which can be viewed in the same spirit as the Hegselmann-Krause opinion dynamics over infinitely many agents under a simplified neighborhood structure. To our knowledge, no work in the data clustering literature directly considers an infinite data set. There are, however, some recent works in the probability literature that address problems similar to ours. Baccelli and Khaniha [khaniha2025hierarchical] consider an infinite-dimensional clustering model on $\mathds{R}^{2}$; however, they are unable to establish a scaling limit for their dynamics. Angel, Ray, and Spinka [angel2023tale] study a model on $\mathds{R}$ that behaves similarly to clustering, although it is not strictly a clustering model. They leave it as an open problem to determine whether a scaling limit exists for their dynamics. Even identifying the correct scaling is a challenging problem for either of these two papers.
1.2 Contributions and Organization of this Paper
In this paper, we prove that there is a scaling limit for Algorithm 1. That is, (under mild assumptions on the initial data), repeated application of Algorithm 1 converges to a unique limit which is independent of the initial data.
@@ 1.2 Contributions and Organization of this Paper
This paper is organized as follows. In Section 2, we describe the model and main results. In Section 3, we prove the main results. In Section 4, we state a conjecture which generalizes our main results. In Section 5, we conclude and discuss future work.
@@ 2.1 Point Process Model
Let $\overrightarrow{\Xi}^{(0)}\in\mathbf{\Xi}$. We define first a point process $\hat{\Xi}^{(t+1)}$ in terms of $\overrightarrow{\Xi}^{(t)}$ as follows. Each point moves halfway towards one of its neighbors, chosen uniformly at random and independently of other points’ movements. If two points move to the same location, they merge so that $\hat{\Xi}^{(t+1)}$ is a simple stationary point process with intensity $\frac{3}{4}$. Then,
@@
is obtained by re-scaling $\hat{\Xi}^{(t+1)}$ to unit intensity. Here, we denote the re-scaling to unit intensity by left-multiplying a point process by its intensity: that is, if $\Xi=(\Xi_{i})_{i\in\mathds{Z}}$, then $c\times\Xi:=(c\Xi_{i})_{i\in\mathds{Z}}$, for any constant $c>0$.
Figure 2 shows our dynamics, without the re-scaling to unit-intensity:
@@ 2.2 Gap Sequence Model
We consider a second formulation of the model, in which the dynamics can be considered as a product of certain $\mathds{Z}$–indexed random matrices. In this formulation of the model, rather than track the point locations, we track the gaps between points. The original point process model can be recovered by also tracking a real-valued quantity representing the location of the 0–indexed point.
@@
We define two random linear operators, $\mathcal{A}^{(t)}$ and $\mathcal{F}^{(t)}$ as follows, which refer to as averaging and folding, respectively. One iteration of the dynamics is expressed by the dynamics
@@
where averaging corresponds to the point movements, and folding corresponds to merging of co-located points. We express these operators as $\mathds{Z}^{2}$–indexed random matrices. To do this, we use the notation $e_{i}$ to represent the $\mathds{Z}$–indexed standard basis vector equal to $1$ in the $i$-th component and equal to 0 in all other components. We also use the notation $\mathcal{A}^{(t)}_{i},\mathcal{F}^{(t)}_{i}$ to denote the $i$-th rows of $\mathcal{A}^{(t)},\mathcal{F}^{(t)}$ respectively, where $i\in\mathds{Z}$. Note that, we have for any $t\geqslant 1$,
$\overrightarrow{\Gamma}^{(t)}=\mathcal{F}^{(t-1)}\mathcal{A}^{(t-1)}\cdots\mathcal{F}^{(2)}\mathcal{A}^{(2)}\mathcal{F}^{(1)}\mathcal{A}^{(1)}\overrightarrow{\Gamma}^{(0)},$ (2.1)
which implies that for any $i\in\mathds{Z}$, the element $\overrightarrow{\Gamma}^{(t)}_{i}$ can be written as a non-negative random linear combination of the entries in $\overrightarrow{\Gamma}^{(0)}$.
@@
$\mathcal{A}^{(t)}_{i}:=\begin{cases}\frac{1}{2}(e_{i}+e_{i+1})&\text{ with probability }1/2\\ \frac{1}{2}(e_{i}+e_{i-1})&\text{ with probability }1/2.\end{cases}$
@@
Folding: Observe that point merges at time $t$ occur according to a stationary $\mathds{Z}$–valued renewal process $\tau^{(t)}$, and that for $t\neq s$, $\tau^{(t)}$ and $\tau^{(s)}$ are i.i.d.. It is easy to see that the inter-renewal distribution of $\tau^{(t)}$ is the sum of two independent $\operatorname{Geom}(1/2)$ random variables. From here, it follows that the rows of the folding matrix $\mathcal{F}^{(t)}$ can be expressed as follows. For a process $\tau\subseteq\mathds{Z}$, let $N_{\tau}(\cdot)$ denote the counting distribution of $\tau$, where
@@
$\mathcal{F}^{(t)}_{i}:=\begin{cases}\frac{3}{4}(e_{i+N_{\tau}(i)-1}+e_{i+N_{\tau}(i)})&\text{ for }i\in\tau^{(t)}\cap[0,\infty)\\ \frac{3}{4}(e_{i+N_{\tau}(i)}+e_{i+N_{\tau}(i)+1})&\text{ for }i\in\tau^{(t)}\cap(-\infty,0)\\ \frac{3}{4}e_{i+N_{\tau}(i)}&\text{ for }i\not\in\tau^{(t)}.\end{cases}$
At the indices of the mergings, we need to correspondingly add the two merged gaps; otherwise, we keep the original gap. In order to do this, we also need to keep track of the number of mergings over space.

### v2

@@ Abstract
We consider an infinite-dimensional stochastic clustering model on $\mathds{R}$. In discrete time, each point of a unit-intensity simple point process moves halfway toward either of its left or right neighbors, chosen uniformly at random. Co-located points are merged into a single point, and the resulting simple point process is rescaled to unit intensity. We show that, when the point processes are shifted so that there is a point at the origin, the dynamics have a unique weak limit when the initial point process is renewal. For this limiting point process, the gap distribution has exponential tails. We also show that for the time-reversed process and with an appropriate scaling in space, there is a limiting (random) distribution function on $\mathds{R}$, whose associated measure assigns to $\mathds{R}$ a measure corresponding to the gap between consecutive points. Finally, we discuss several relevant research directions.
@@ 1 Introduction
Dynamic clustering refers to a class of stochastic or deterministic dynamics in which a configuration of data points (or particles) evolves over time through local interactions [2]. These interactions cause points to aggregate, merge, split, or reorganize, so that the resulting cluster structure is itself time-dependent and emerges from the underlying dynamics. Dynamic clustering is often done via numerical optimization [29]. For large, finite datasets, there are two challenges with this approach. First, large numerical optimization problems are computationally expensive. Second, and perhaps more important, is the question of determining stopping criteria for the clustering algorithm. For finite datasets, continued use of a clustering algorithm often results in all data points being assigned to a single cluster, which is undesirable.
In this work, we address this second question. Specifically, we study stochastic dynamic clustering algorithms on infinite datasets, aiming to determine whether these dynamics possess a stationary measure. If a stationary measure exists, a natural stopping criterion for the clustering dynamics is when one is close to that stationary measure. For large finite datasets, one can also accept the current clustering when the cluster gap distribution is close to that of the stationary measure for a corresponding infinite dataset.
To our knowledge, this is the first work to directly analyze this type of clustering dynamics on an infinite dataset, with the intent of informing behavior on large finite datasets. Indeed, we demonstrate this behavior via simulation for two simple algorithms below. The algorithms are as follows.
@@
In this paper, we prove that Algorithm 1 has a scaling limit. That is, (under mild assumptions on the initial data), repeated application of Algorithm 1 converges to a unique limit that is independent of the initial data. Moreover, the size of a typical cluster (the number of points merged into a single point) after appropriate scaling converges to a nontrivial random variable. Finally, we prove the existence of a random measure whose mass gives the limiting scaled gap distribution, and whose length of support gives the limiting scaled cluster size distribution.
Note that while Algorithms 1 and 2 seem similar, even identifying the scaling required in Algorithm 2 is a nontrivial task. For Algorithm 1, the factor $\frac{3}{4}$ appears due to the fact that neighboring points will co-locate with probability $\frac{1}{4}$, independently of their spatial positions. It is far from obvious what the correct scaling factor should be for Algorithm 2, and even if that scaling factor should be independent of the initial data. Thus, there is a significant amount of interesting and important future work to be done in terms of understanding the limiting properties of these kinds of dynamics.
@@ 1.1 Related Work
Clustering problems on static data have been widely studied in various areas, such as the minimum distortion problem in data compression [22], facility location problems [18], and $k$-means clustering in pattern recognition [30], among many others [37]. However, recent advances in geographic information systems have led to growing interest in developing dynamic clustering algorithms for ensembles of moving objects or for data points that evolve over time. Applications of dynamic clustering include coverage control for mobile sensing networks [13, 33, 38], the development of automatic deployment and tracking/monitoring algorithms for surveillance systems [20, 39], modeling database demand variability and data placement in web applications [23, 19], clustering spatio-temporal brain signal dynamics [32], and traffic routing through traffic-flow clustering [34].
From a different perspective, dynamic clustering has also emerged in a wide range of multi-agent averaging dynamics, which have been studied extensively from a control-theoretic viewpoint, with applications ranging from distributed coordination and formation [25, 26], to modeling animal flocking patterns [12], and to opinion dynamics [14, 21, 24]. While averaging dynamics were initially used to model consensus in multi-agent systems (e.g., robotic rendezvous [10] or consensus in social networks [14]), these models were later extended to study disagreement (clustering) in addition to consensus.
@@
For instance, in the context of opinion dynamics, Friedkin and Johnsen were the first to adopt this approach, proposing a dynamical system that captures opinion polarization into multiple clusters [21]. This model was later extended by Hegselmann and Krause [24] and belongs to the broader class of bounded-confidence dynamics [6]. Beyond bounded-confidence models, dynamic clustering has also been studied in broader classes of nonlinear and signed interaction systems, where antagonistic or repulsive interactions naturally lead to the emergence of multiple opinion groups [4, 6]. Analyzing the convergence and limit points of such dynamics, however, poses major challenges compared to standard averaging models, whose convergence properties can often be studied using classical tools from Markov chains or graph theory. These difficulties become even more pronounced when additional stochasticity is introduced into the dynamics [36, 16].
@@
As another example, averaging dynamics on state-dependent graphs (e.g., where interactions depend on agents’ positions) are closely related to dynamic clustering, especially when the data points are spatially distributed and evolve over time, and the goal is to cluster them based on similarity [24, 1, 7]. Similarly, in dynamic community detection, one aims to understand how diffusion-based averaging dynamics on time-varying graphs can reveal communities (clusters) that persist over time [15]. In both cases, the co-evolution of agents’ states and network structure introduces dynamic coupling among agents, requiring new analytical approaches to characterize limit points and long-term behavior. These complementary perspectives highlight that dynamic clustering is a robust and structurally rich outcome of state-dependent interactions, extending well beyond classical consensus frameworks.
There are a few references on averaging processes on graphs [3], in which an edge is selected at random, and the incident vertex values are averaged. This model has been generalized to hypergraphs in [35]. Recent work by [11] studies the cutoff phenomenon for the average process on the complete graph, in connection with a question by Bourgain related to quantum computing. However, our model differs from the average process, as all particles move simultaneously at each time step. In Section 3.6, we give an example of a toy model of follow-the-leader that exhibits behavior similar to the average process and provides motivation for analyzing the dual weight process and its distribution function.
@@
Motivated by these applications, we propose and study a new class of averaging dynamics for spatial data clustering, which can be viewed in the same spirit as the Hegselmann-Krause opinion dynamics [24], but over infinitely many agents and under a simplified neighborhood structure. To our knowledge, no work in the data clustering literature directly considers an infinite data set. There are, however, some recent works in the probability literature that address problems similar to ours. Baccelli and Khaniha [28] consider an infinite-dimensional clustering model on $\mathds{R}^{2}$; however, they are unable to establish a scaling limit for their dynamics. Angel, Ray, and Spinka [5] study a model on $\mathds{R}$ that behaves similarly to clustering, although it is not strictly a clustering model. They leave the question of whether a scaling limit exists for their dynamics as an open problem. Even identifying the correct scaling is a challenging problem for either of these two papers. Further, identifying the scaling for Algorithm 2 is also a challenging problem; our results are facilitated by the straightforward scaling for Algorithm 1.
1.2 Contributions and Organization
The main contributions of this work are as follows.
  • We propose and analyze a new class of dynamical clustering algorithms for spatial $\mathbb{R}$-valued data, where each point randomly clusters with its left or right neighbor. To our knowledge, this is the first study to examine such dynamics on an infinite dataset.
  • For this model, we show that if the initial point process is a renewal process and the point processes are shifted so that a point lies at the origin, a unique weak limit exists for both scaled positions and cluster sizes, which is independent of the initial data.
  • Our proof combines a time-reversal construction with stochastic duality and a recursive analysis of reverse-time weights, which may be useful for other infinite-dimensional clustering dynamics. The results about concentration for stationary renewal processes may also be of independent interest.
Our analysis of Algorithm $1$ relies crucially on the following two properties. First, we have order preservation: that is, if a point $u$ starts to the left of a point $v$, it will never be to the right of $v$ (this property also holds for Algorithm 2). Second, the time-reversal dynamics can be described as the composition of certain independent (random) linear operators. This provides a well-behaved structure from which we can show our results. Independence in the second property fails for Algorithm 2.
@@ 1.2 Contributions and Organization
This paper is organized as follows. In Section 2, we provide relevant background materials and notations. In Section 3, we describe the model and main results. We prove the main results in Section 4. In Section 5, we prove the technical results in this article. Finally, in Section 6, we discuss future work.
2 Background Material and Notations
In this section, we first provide definitions and preliminary results, along with the notations used throughout the paper.
Definition 2.1.
[…]
Definition 2.2.
A point process $\Xi$ on $\mathds{R}$ is non-simple if
[…]
For a non-simple point process $\Xi$, the simplified point process corresponding to $\Xi$ is a simple point process with exactly one point at each location where $\Xi$ has a point.
Definition 2.3.
A point process $\Xi$ is stationary if for any two sets $S_{1},S_{2}\subset\mathds{R}$ of equal and finite Lebesgue measure, the distributions of the number of points in $S_{1}$ and $S_{2}$ are equal. If $\Xi$ is a stationary point process, the Palm-shifted point process $\Theta\Xi$ is the point process constructed by shifting all points of $\Xi$ such that the 0-indexed point of $\Xi$ is at the origin.
These definitions can be readily restricted to define point processes on $\mathbb{Z}$, and we will do so without further mention.
Definition 2.4.
Let $N(t)$ denote the number of points of a stationary point process $\Xi$ in $(0,t]$, for $0\neq t\in\mathds{R}$. The intensity of $\Xi$ is $\lambda:=\lim_{t\to\pm\infty}\operatorname{\mathds{E}}N(t)/t$, provided it exists.
The intensity of a point process can be readily extended to be spatially inhomogeneous or even random, but we will not require that level of generality in this paper; thus, we provide only this restricted definition.
Lemma 2.5 ([27, Theorem 6.1]).
For Palm-shifted point processes, convergence of finite-dimensional distributions implies tightness. Thus, convergence of finite-dimensional distributions is equivalent to weak convergence.
Definition 2.6.
Let $\Gamma^{(t)}$ and $\eta^{(t)}$ be Markov processes on measurable spaces $S_{1}$ and $S_{2}$, respectively. The processes $\Gamma^{(\cdot)}$ and $\eta^{(\cdot)}$ are stochastically dual with respect to a function $h:S_{1}\times S_{2}\to\mathds{R}$ if for all $t\in\mathds{Z},\Gamma^{(0)}\in S_{1},\eta^{(0)}\in S_{2}$, we have:
$\operatorname{\mathds{E}}_{\Gamma^{(0)}}h(\Gamma^{(t)},\eta^{(0)})=\operatorname{\mathds{E}}_{\eta^{(0)}}h(\Gamma^{(0)},\eta^{(t)}).$
The intuitive idea of stochastic duality is that if certain functionals of a Markov process are difficult to compute, those same functionals may be more readily computed for a stochastic dual process. In this paper, we use time reversal as the stochastic dual.
2.1 Notations
We adopt the following notation in this paper. For a process $X$, $\overrightarrow{X}^{(t)}$ denotes the forward-time evolution, and $\overleftarrow{X}^{(t)}$ denotes the reverse-time evolution. For a point process $\Xi$, $\Theta\Xi$ denotes the Palm-shifted point process. Convergence of random variables will be denoted as follows: $X_{n}\xrightarrow{\mathrm{f.d.d.}}X$ denotes convergence of finite-dimensional distributions; $X_{n}\xrightarrow{L^{p}}X$ denotes convergence in $L^{p}$; and $X_{n}\Rightarrow X$ denotes weak convergence.
@@ 3.1 Point Process Model
Let $\overrightarrow{\Xi}^{(0)}\in\mathbf{\Xi}$. We first define a point process $\hat{\Xi}^{(t+1)}$ in terms of $\overrightarrow{\Xi}^{(t)}$ as follows. Each point moves halfway towards one of its neighbors, chosen uniformly at random and independently of other points’ movements. If two points move to the same location, they merge so that $\hat{\Xi}^{(t+1)}$ is a simple stationary point process with intensity $\frac{3}{4}$. Then,
@@
is obtained by re-scaling $\hat{\Xi}^{(t+1)}$ to unit intensity. Here, we denote the re-scaling to unit intensity by left-multiplying a point process by its intensity: that is, if $\Xi=(\Xi_{i})_{i\in\mathds{Z}}$, then $c\times\Xi:=(c\Xi_{i})_{i\in\mathds{Z}}$, for any constant $c>0$. We refer to Figure 2 for an illustration of our dynamics without rescaling to unit intensity.
We remark that we could instead define the dynamics on the space of non-simple point processes. In that case, our dynamics preserve the unit-intensity trivially. However, the resultant non-simple point processes consist of ever-increasing numbers of co-located points at ever-increasing distances, and it is difficult to make sense of a stationary measure. It is more convenient to scale the simplified point process and endow each point with an $\mathds{N}$-valued mark equal to the number of initial points co-located at that point in the simplified point process.
@@
3.1.1 Informal Statements of Results
Here, we state our main results informally, followed by their formal statements in Section 3.4.
First, we prove that the (Palm-shifted) dynamics of $\Xi^{(t)}$ are ergodic; that is, they admit a unique limit point. We then show that the tail of the limiting gap distribution is exponential. We also establish a scaling limit for the number of initial points that have merged with the initially 0-indexed point. These results are obtained via a time-reversal argument; in addition, we show that the time-reversed process is a stochastic dual of the original dynamics.
Our result is about the ergodicity of a certain infinite-dimensional Markov process that describes Algorithm 1. For dynamic clustering, this result means that independently of the initial data, the distribution near which the clustering should be accepted is the same. However, the limiting (non-simple before cluster size scaling) point process is not renewal: there are dependencies between the size and locations of neighboring clusters that arise. Thus, the behavior of Algorithm 1 is a “smoothing” of the initial data in some sense, but nevertheless it is this “smoothing” that allows us to prove the result – for infinite dimensional Markov processes, it is not typical for a stationary measure to be unique. Notice that in Figure 1, this behavior does not seem to hold for Algorithm 2. This is a more “natural” behavior, but it requires the development of fundamentally different tools than the ones we employ here for the simpler Algorithm 1.
In Theorem 3.5, we examine the dynamics of the time-reversed process in greater detail to show how the limiting gap distribution can be iteratively constructed; this requires the gap sequence model and its time reversal as defined in Section 3.3.
Remark 1.
In our model, each point moves towards its right neighbor independently with probability $\frac{1}{2}$, so the scaling factor is $\frac{3}{4}=1-\frac{1}{2}\cdot\frac{1}{2}$. For a model where points move to the right with probability $p$ and to the left with probability 1p1-p, one can study it similarly with the scaling 1p(1p)1-p(1-p); however, there will be drift at speed 2p12p-1. This drift is obviously to the right when $p>\frac{1}{2}$ and to the left when $p<\frac{1}{2}$, and that can be interpreted from the sign of 2p12p-1. All of our results have analogues in this more general situation.
We expect that analogues of our results also hold when each particle moves halfway towards one of its $k$ nearest neighbors in each direction. In this setting, we no longer have order conservation: a point $u$ that starts to the left of another point $v$ could end up to the right of $v$. This complicates the analysis significantly. In the model we study, a point $u$ that starts at another point $v$’s left either remains to the left of $v$ or co-locates with $v$.
@@ 3.2 Gap Sequence Model
We consider a second formulation of the model, in which the dynamics can be considered as a product of certain $\mathds{Z}$–indexed random matrices. In this formulation of the model, rather than tracking the point locations, we track the gaps between points. The original point process model can be recovered by also tracking a real-valued quantity representing the location of the 0–indexed point.
@@
We define two random linear operators, $\mathcal{A}^{(t)}$ and $\mathcal{F}^{(t)}$, which we refer to as averaging and folding, respectively. One iteration of the dynamics is expressed by the dynamics
@@
where averaging corresponds to the point movements, and folding corresponds to merging of co-located points.
To define these operators formally as $\mathbb{Z}^{2}$-indexed random matrices, let us use the notation $\mathcal{A}^{(t)}_{i},\mathcal{F}^{(t)}_{i}$ to denote the $i$-th rows of $\mathcal{A}^{(t)},\mathcal{F}^{(t)}$ respectively, where $i\in\mathds{Z}$. Note that we have,
$\overrightarrow{\Gamma}^{(t)}=\mathcal{F}^{(t-1)}\mathcal{A}^{(t-1)}\cdots\mathcal{F}^{(2)}\mathcal{A}^{(2)}\mathcal{F}^{(1)}\mathcal{A}^{(1)}\overrightarrow{\Gamma}^{(0)}\ \ \ \forall t\geqslant 1,$ (3.1)
which implies that for any $i\in\mathds{Z}$, the element $\overrightarrow{\Gamma}^{(t)}_{i}$ can be written as a non-negative random linear combination of the entries in $\overrightarrow{\Gamma}^{(0)}$. Also, let $e_{i}$ denote the $\mathbb{Z}$-indexed standard basis vector that is equal to $1$ in the $i$-th component and 0 in all other components.
@@
$\mathcal{A}^{(t)}_{i}:=\begin{cases}\frac{1}{2}(e_{i}+e_{i+1})&\text{ with probability }1/2,\\ \frac{1}{2}(e_{i}+e_{i-1})&\text{ with probability }1/2.\end{cases}$
@@
Folding: Observe that point merges at time $t$ occur according to a stationary $\mathds{Z}$–valued renewal process $\tau^{(t)}$, and that for $t\neq s$, $\tau^{(t)}$ and $\tau^{(s)}$ are i.i.d. It is easy to see that the inter-renewal distribution of $\tau^{(t)}$ is the sum of two independent $\operatorname{Geom}(1/2)$ random variables. From here, it follows that the rows of the folding matrix $\mathcal{F}^{(t)}$ can be expressed as follows. For a process $\tau\subseteq\mathds{Z}$, let $N_{\tau}(\cdot)$ denote the counting function of $\tau$, where
@@
$\mathcal{F}^{(t)}_{i}:=\begin{cases}\frac{3}{4}(e_{i+N_{\tau}(i)-1}+e_{i+N_{\tau}(i)})&\text{ for }i\in\tau^{(t)}\cap[0,\infty),\\ \frac{3}{4}(e_{i+N_{\tau}(i)}+e_{i+N_{\tau}(i)+1})&\text{ for }i\in\tau^{(t)}\cap(-\infty,0),\\ \frac{3}{4}e_{i+N_{\tau}(i)}&\text{ for }i\not\in\tau^{(t)}.\end{cases}$
At the indices of the mergings, we need to correspondingly add the two merged gaps; otherwise, we keep the original gap. To do this, we also need to keep track of the number of mergings across space.